# Fast Real Evaluation Through Sound Mixed-Precision Tuning

### v1

@@ 7.2. Results
Besides the core research questions discussed here, Reval’s precision assignment algorithm provides several additional advantages. One is robustness: on the approximately 72,450 points that none of the three tools could evaluate, Reval terminates much faster. This is because Reval can exit after assigning very high precisions, which takes only one or two iterations. For these points, Reval terminates 11.09×11.09$\times$11.09 × faster than Sollya and 19.87×19.87$\times$19.87 × faster than Baseline. While these points are not considered in RQ1–RQ3, they play a large role in practical deployments of real evaluation, to the extent that the underlying Rival library uses special flags to detect some kinds of unevaluable inputs (Flatt and Panchekha, 2023). Figure 5 breaks down Reval’s greater advantage once unevaluable points are included: early exit allows Reval to skip many useless and expensive high-precision operations, while the uniform precision Baseline has to execute these operations to reach the maximum precision. This also makes it practical for Reval to use a significantly higher maximum precision in practical deployments, which in turn allows it to evaluate more inputs. Finally, compared to Sollya, Reval supports a broader range of operations (cbrt and atan2 are two examples which show up in the Herbie 2.0 benchmark suite), produces correctly- not faithfully-rounded results, and never returns incorrect results, Sollya produces faithfully-, but not correctly-rounded results in 5,000 cases.
Figure 5. A variant of Figure 4, but over all inputs, including those that neither Reval nor Baseline could evaluate. Reval’s early exit allows it to skip most high-precision evaluations, while Baseline wastes a substantial fraction of its runtime uselessly re-evaluating expressions at ever-higher precisions.
@@ 7.3. Case Studies
We now examine individual expressions and input points in more detail to highlight aspects of Reval’s precision tuning algorithm.
@@
The 2cos benchmark evaluates $f(x,\varepsilon)=\cos(x)-\cos(x+\varepsilon)$ to 53 mantissa bits. When $x$ is big and $\varepsilon$ is small, for example $x=10^{300}$ and $\varepsilon=10^{-300}$, the terms being subtracted differ by a tiny amount, making the output hard to compute. To evaluate the output correctly, the $x+\varepsilon$ term must be computed to at least 600 digits, or about 2000 bits. Then the $\cos$ operations have to be evaluated to approximately 300 digits, or about 1000 bits, and finally the subtraction needs to be evaluated to only 53 bits to achieve a correctly-rounded result.
Reval achieves this precision assignment in three iterations. In the first, untuned iteration, Reval is able to estimate the sizes of $x$, $x+\varepsilon$, and $\cos(x)$. However, because this first iteration uses low precision, $x+e$ covers more than one full period of $\cos$—in fact, it covers approximately $3\cdot 10^{282}$ of them. The interval computed for $\cos(x+e)$ is thus $[-1,1]$, and the subtraction operation likewise has a wide computed interval crossing zero. As a result, during the first tuning pass, Reval applies the slack mechanism twice, once to the subtraction and once to cosine. This results in a precision assignment of $59073+5+\operatorname{slack}[0]=$ 590 for both $\cos$ operations, the slack being introduced by the subtraction, and $590+5+1000+\operatorname{slack}[0]\approx 2100$ for $x+\varepsilon$, where $590$ is the output precision needed, $1000$ is the amplification factor for cosine, and the slack being introduced by the cosine. The amplification factor of $1000$ is a good demonstration of how slack differs from naive uniform precisions. Even though the slack mechanism is used, the $\cos$ operation still provides some information. Specifically, the input to $\cos$ has to be precise enough to at least identify what period of $\cos$ it is in. The full precision assignment is shown in Figure 6.
The first Reval precision assignment:
$\begin{array}[]{llll}P[-]&&&=73\\ P[\cos(x)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil&=73+5+\operatorname{slack}[0]&=590\\ P[\cos(+)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{2}(-)\rceil&=73+5+\operatorname{slack}[0]&=590\\ P[+]&=P[\cos(+)]+5+\lceil\log_{2}\operatorname{ampl}_{1}(\cos(+))\rceil&=590+5+1000+\operatorname{slack}[0]&=2107\end{array}$
The second Reval precision assignment:
$\begin{array}[]{llll}P[-]&&&=73\\ P[\cos(x)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil&=73+5+593+\operatorname{slack}[1]&=1695\\ P[\cos(+)]&=P[-]+5+\lceil\log_{2}\operatorname{ampl}_{2}(-)\rceil&=73+5+593+\operatorname{slack}[1]&=1695\\ P[+]&=P[\cos(+)]+5+\lceil\log_{2}\operatorname{ampl}_{1}(\cos(+))\rceil&=1695+5+1000&=2700\\ \end{array}$
@@
Figure 6. Precision assignment in the first and second iteration for the slack mechanism case study. Note that, even though the slack mechanism partially re-introduces precision doubling, amplification factors still provide useful information.
@@
After re-evaluating with these new precision assignments, the output of the $\cos$ operation no longer crosses zero, but the output of the subtraction still does. The second precision assignment, then, uses the slack mechanism only for the subtraction, assigning a precision of $169573+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil=78+593+\operatorname{slack}[1]=$ 1695 to both $\cos$ operations and a precision of $1695+5+1000=2700$ to the $x+\varepsilon$ term. A third re-evaluation then results in a correctly-rounded result. Note that, while Reval needed only 3 evaluations, a uniform-precision evaluation would require 5 iterations.
@@
The placeG benchmark evaluates $f(x,y,z)=(x+y)\times(z+1)$ to 53 mantissa bits. Since the benchmark uses only multiplications and divisions, the exact real result can be represented in floating-point. Consider positive $x$, $y$, and $z$; then all the $\operatorname{ampl}$ and $\operatorname{intro}$ factors are 1 or less. Reval thus assigns low precisions to all operations: 73 bits to the multiplication and 78 to both additions. That works for many positive inputs, but not all.
Some specific inputs suffer from double rounding, such as $x=1.3002052657264033\times 10^{189}$, $y=3.084776002356433\times 10^{188}$, and $z=1\times 10^{-200}$ in double-precision floating-point. These $x$ and $y$ values are exactly one exponent apart, and both end in a 1 bit. This means that their sum, $x+y$, is exactly 54 bits wide, meaning it lies exactly on a rounding boundary. That means it rounds up if $1+z$ is greater than $1$, and down (toward even) if $1+z$ is equal to $1$. In fact, $z+1$ is very slightly greater than 1, which means the answer should round up, but because $z$ is much smaller than $1$, $1+z$ needs to be computed to many, many bits to conclude that it is in fact greater than $1$.
Specifically, in 78 bits, $z+1$ is computed to be $[1,1.000000000000000000000007]$ in the first iteration and the overall expression is computed to be $[1.6...\times 10^{189},1.6...\times 10^{189}]$, where the two endpoints of that interval are successive floating-point numbers. This triggers the double rounding mechanism, causing Reval to evaluate the expression first to $58573+\operatorname{slack}[0]=$ 585 and then to 73+slack[$109773+\operatorname{slack}[1]=$ 1097 bits. This last iteration finally evaluates $1+z$ exactly, proving that it is strictly greater than $1$. By increasing the overall precision demands, the double rounding mechanism was able to eventually determine which side of the rounding boundary the result was on.
@@
The 2nthrt benchmark evaluates $f(x,n)=(x+1)^{1/n}-x^{1/n}$ to double precision. The benchmark is hard to evaluate for large $x$ and small $n$—say, $x=10^{200}$ and $n=10^{-200}$. These inputs require raising a large number to a large power, causing overflow: the intermediate result of $(x+1)^{1/n}$ is $[\approx 2\times 10^{323,228,496},+\infty]$ after the first iteration. The $x^{1/n}$ term likewise overflows, resulting in an output interval of [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] after subtracting the two.
When adjusting precisions for the second iteration, precision tuning algorithm calculates an amplification factor for the subtraction of
@@
where each term contributes a slack term. This causes Reval to assign a precision of billions of bits to the left-hand term, higher than the maximum allowed precision 32 256. Reval therefore terminates evaluation early.
@@ 8. Related Work
H. Boehm was an early advocate of evaluating real-number expressions to high precision using “constructive reals” (Boehm, 1987). A number of implementations of constructive reals exist, including in Java (Boehm, 2004), Python (Briggs, 2006), Caml (Ménissier-Morain, 2005), and others. Perhaps most widely used is the implemenation (Boehm, 2020) in the Android Calculator application. Early implementations used lazy digit sequences, but later work (Lee and Boehm, 1990; Boehm et al., 1986) showed that Ziv’s strategy (Ziv, 1991) with interval arithmetic is typically faster. Interval arithmetic is an old and widely-used technique. Prominent implementations include Arb (Johansson, 2016), Moore (Mascarenhas, 2016), MPFI (Revol and Rouillier, 2005), Rival (Flatt and Panchekha, 2023), and others. Interval arithmetic is used in the Sollya environment (Chevillard et al., 2010) for numerical programming and possibly in Wolfram Mathematica (Flatt and Panchekha, 2023). Many researchers have proposed various tweaks to computable real implementations, such as movability flags to track overflow (Flatt and Panchekha, 2023). Perhaps the most similar to the current work was proposed by Li and Yong (2007): mixed-precision evaluation using “weights”. However, they seem to expect the weights to be user-controlled, whereas Reval derives the precision tuning automatically.
Interval arithmetic and constructive reals are closely related to research into automatic worst-case error bounds for floating-point programs. One important early system was Salsa (Damouche and Martel, 2017; Martel, 2009; Damouche et al., 2015), which used interval arithmetic to bound both ranges of variables and also errors, a strategy adopted by other work as well (Titolo et al., 2018). Rosa (Darulova and Kuncak, 2014) and Daisy (Izycheva and Darulova, 2017) use a similar analysis except with affine arithmetic, which better handles correlated errors. Initially phrased as an abstract interpretation, later papers (Martel, 2018) rephrased the analysis as a type system, where types track both range and precision. FPTaylor (Solovyev et al., 2018) introduced the idea of error Taylor series for automatic worst case error bound estimation, achieving much lower error. Satire (Das et al., 2020) then proposed scaling error Taylor series using abstraction, algebraic simplification, and automatic differentiation, and SeeSaw (Das et al., 2021) extended these techniques to handle bounded control flow. Like Reval, Satire saw higher-order error as a computational bottleneck; unlike Reval, Satire sacrificed soundness for performance, while Reval instead accepted larger error bounds. ATOMU (Zou et al., 2019) proposed computing first-order error from condition numbers. While this provides an algorithmic speedup, ATOMU did not evaluate it; instead, ATOMU focused on condition numbers as a method for driving input generation. In most worst-case error bound tools, the goal is achieving the lowest possible sound error bound, unlike in this paper, where the focus is primarily on computational efficiency.
@@ 9. Conclusion
Evaluating real expressions to high precision requires using high-precision interval arithmetic. But state-of-the-art tools use a uniform precision, doubled with every iteration. This wastefully allocates too much precision to non-accuracy-critical operations, and can also overestimate precision or require too many iterations to find the right precision. We instead introduce a fast, sound method for assigning precisions to interval operators so as to correctly-round a real expression. Our method is based on a variation on error Taylor series and condition numbers, called amplification factors, which is sound, computable quickly, and specialized to interval arithmetic. A key optimization is the use of “exponent tricks”, which allow computing amplification factors using only operations on machine integers. As a result, Reval is able to evaluate real expressions 1.25×1.25$\times$1.25 × faster than the state-of-the-art Sollya tool across our suite of 528 benchmarks, with Reval’s advantage rising to 2.99×2.$9$ for the most difficult input points. It achieves this speedup largely by using a sophisticated mixed-precision assignment algorithm that evaluates operations at high precision only when necessary.

### v2

@@ 7.2. Results
One additional advantage of Reval is its handling of unevaluable inputs, where it (and Baseline and Sollya) could not evaluate the input due to the maximum precision. For these inputs, Reval terminates 6.58×6.58$\times$6.58 × faster than Baseline (and 13.32×13.32$\times$13.32 × faster than Sollya). Reval is much faster than Baseline on these inputs because it can typically assign a very high precision after just one or two iterations and then perform an early exit, while Baseline must try every uniform precision before giving up. Figure 7 replicates Figure 3 but with these unevaluable inputs included: Baseline now spends a huge amount of time on fruitless evaluations at high precision, while Reval spends very little additional time. While these inputs are not considered in RQ1–RQ3, they play a large role in practical deployments, as evidenced by prior work on movability flags to detect some such inputs (Flatt and Panchekha, 2023). Because unevaluable inputs can now be swiftly dispatched, non-uniform precision assignments also allow setting a much higher maximum precision; with a uniform precision, raising the maximum precision dramatically raises the cost of unevaluable inputs.
Figure 7. A variant of Figure 3, but over all inputs, including those that neither Reval nor Baseline could evaluate. Reval’s early exit optimizations allow it to skip high-precision evaluations of unevaluable inputs, while Baseline must waste time performing. This makes Reval more robust to unevaluable inputs and could enable users to set higher maximum precisions.
Early exit does have some costs: for 47 inputs (of the 567,238 total), Baseline produces a correctly-rounded result while Reval does not. This is typically because Reval assigns a precision just over the maximum precision and exits early, while Baseline tries the maximum precision and succeeds. This extremely rare occurence can be prevented by enabling Reval’s “lower bound” mode; this allows Reval to evaluate all points that Baseline does, but adds a 0.98% slowdown to Reval. Note that this slowdown figure is somewhat deceptive, since it excludes unevaluable points, on which it forces a re-evaluation, at maximum precision, of 18,057 points, causing a significant slow-down. As a result, we make this “lower bound” mode off by default: it addresses a very rare issue and hurts robustness to unevaluable inputs. Ultimately, we expect Reval users to choose much higher maximum precisions than they would with a uniform precision due to the robustness advantages noted above.
@@ 7.3. Case Studies
We now illustrate three more-complex features of Reval—the slack mechanism, correct rounding, and early exit—by case studies on individual inputs.
@@
$\begin{array}[]{llllll}z_{i}&\text{First tuning iteration}&\text{Second tuningiteration}\\ \hline\cr z_{4}=z_{3}-z_{2}&53+10=63&53+10=63\\ z_{3}=\cos(x)&63+5+\operatorname{slack}[0]=580&63+5+(593+\operatorname{slack}[1])=1685\\ z_{2}=\cos(z_{1})&63+5+\operatorname{slack}[0]=580&63+5+(593+\operatorname{slack}[1])=1685\\ z_{1}=x+\varepsilon&580+5+(1000+\operatorname{slack}[0])=2097&1685+5+1000=2690\\ \end{array}$
@@
Figure 8. Precision assignment in the first and second iteration for the slack mechanism case study. Note that, even though the slack mechanism partially re-introduces search, amplification factors still provide useful information and Reval still converges to a high precision much more quickly than Baseline.
Consider the 2cos benchmark, $\cos(x)-\cos(x+\varepsilon)$, with inputs $x=10^{300}$ and $\varepsilon=10^{-300}$. [^22] This input (and others in the case studies) are a simplified form of one of the 567,238 inputs. The terms being subtracted differ by a tiny amount, and need to be evaluated to approximately 300 digits (1000 bits) to achieve a correctly-rounded result. This in turn requires evaluating $x+\varepsilon$ to at least 600 digits, or about 2000 bits. Reval achieves this precision assignment in two tuning iterations, shown in Figure 8.
After the first, untuned evaluation, Reval is able to estimate the sizes of $x$, $x+\varepsilon$, and $\cos(x)$. However, because this first iteration uses low precision, $x+e$ covers a wide interval and the interval computed for $\cos(x+e)$ is thus $[-1,1]$, and the subtraction likewise has a wide interval crossing zero. As a result, during the first tuning pass, Reval applies the slack mechanism twice, once to the subtraction and once to cosine, ultimately assigning a precision of $58063+5+\operatorname{slack}[0]=$ 580 for both $\cos$ operations, the slack being introduced by the subtraction, and $2097580+5+1000+\operatorname{slack}[0]=$ 2097 for $x+\varepsilon$, where $1000$ is the amplification factor for cosine, and the slack is introduced by the cosine. This amplification factor of $1000$ is a good demonstration of how slack differs from naive uniform precisions: even though the slack mechanism is used, the $\cos$ operation still provides some information and allows Reval to make a more useful guess.
@@
After re-evaluating with these new precision assignments, the output of the $\cos$ operation no longer crosses zero, but the output of the subtraction still does. The second precision assignment, then, uses the slack mechanism only for the subtraction, assigning a precision of $168563+5+\lceil\log_{2}\operatorname{ampl}_{1}(-)\rceil=68+593+\operatorname{slack}[1]=$ 1685 to both $\cos$ operations and a precision of $1685+5+1000=2690$ to the $x+\varepsilon$ term. A third re-evaluation then results in a correctly-rounded result. Note that, while Reval needed three evaluations, a uniform-precision evaluation would require evaluations at $2^{6}$, $2^{7}$, $2^{8}$, $2^{9}$, $2^{10}$, and $2^{11}$ before finally succeeding at a precision of $2^{12}$.
@@
Consider the placeG benchmark, $(x+y)\times(z+1)$, at the double-precision inputs $x=1.3002052657264033\times 10^{189}$, $y=3.084776002356433\times 10^{188}$, and $z=1\times 10^{-200}$. Since all inputs are positive and the only operations are addition and multiplication, all $\operatorname{ampl}$ and $\operatorname{intro}$ factors are 1 or less and Reval thus assigns 64 bits of precision to all operations. However, these precisely-chosen $x$ and $y$ values are exactly one exponent apart and both end in a 1 bit. This means that their sum, $x+y$, lies exactly on a rounding boundary, rounding up if $z+1>1$ and down otherwise. This triggers the double rounding mechanism, causing Reval to evaluate the expression first to $57563+\operatorname{slack}[0]=$ 575 and then to 63+slack[$108763+\operatorname{slack}[1]=$ 1087 bits. Only after this second re-evaluation can Reval prove that $z+1$ is strictly greater than $1$ and achieve a correctly-rounded result.
@@
Consider the 2nthrt benchmark, $(x+1)^{1/n}-x^{1/n}$, for $x=10^{200}$ and $n=10^{-200}$. These inputs require raising a large number to a large power, causing overflow: $(x+1)^{1/n}$ evaluates to $[\approx 2\times 10^{323,228,496},+\infty]$ after the first, untuned evaluation. The $x^{1/n}$ term likewise overflows, resulting in an output interval of [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] after subtracting the two.
When adjusting precisions for the second iteration, the amplification factor for the subtraction is
@@
This causes Reval to assign a precision of billions of bits to the left-hand term, higher than the maximum allowed precision 32 256. Reval therefore, correctly, terminates evaluation early.
@@ 8. Related Work
H. Boehm was an early advocate of evaluating real-number expressions to high precision using “constructive reals” (Boehm, 1987). A number of implementations of constructive reals exist, including in Java (Boehm, 2004), Python (Briggs, 2006), Caml (Ménissier-Morain, 2005), and others. Perhaps most widely used is the implemenation (Boehm, 2020) in the Android Calculator application. Early implementations used lazy digit sequences, but later work (Lee and Boehm, 1990; Boehm et al., 1986) showed that Ziv’s strategy (Ziv, 1991) with interval arithmetic is typically faster. Prominent interval arithmetic implementations include Arb (Johansson, 2016), Moore (Mascarenhas, 2016), MPFI (Revol and Rouillier, 2005), Rival (Flatt and Panchekha, 2023), and others. Many researchers have proposed various tweaks to computable real implementations, such as movability flags to track overflow (Flatt and Panchekha, 2023). Perhaps the most similar to the current work was proposed by Li and Yong (2007): mixed-precision evaluation using “weights”. However, this work expects users to manually set weights, whereas Reval derives a non-uniform precision assignment automatically.
Interval arithmetic and constructive reals are closely related to research into automatic worst-case error bounds for floating-point programs. One important early system was Salsa (Damouche and Martel, 2017; Martel, 2009; Damouche et al., 2015), which used interval arithmetic to bound both ranges of variables and also errors, a strategy adopted by other work as well (Titolo et al., 2018). Rosa (Darulova and Kuncak, 2014) and Daisy (Izycheva and Darulova, 2017) use a similar analysis except with affine arithmetic, which better handles correlated errors. Initially phrased as an abstract interpretation, later papers (Martel, 2018) rephrased the analysis as a type system, where types track both range and precision. FPTaylor (Solovyev et al., 2018) introduced the idea of error Taylor series for automatic worst case error bound estimation, achieving much lower error. Satire (Das et al., 2020) then proposed scaling error Taylor series using abstraction, algebraic simplification, and automatic differentiation, and SeeSaw (Das et al., 2021) extended these techniques to handle bounded control flow. Like Reval, Satire saw higher-order error as a computational bottleneck; unlike Reval, Satire sacrificed soundness for performance, while Reval instead accepted larger error bounds. ATOMU (Zou et al., 2019) proposed computing first-order error from condition numbers. While this provides an algorithmic speedup, ATOMU did not evaluate it; instead, ATOMU focused on condition numbers as a method for driving input generation. In most worst-case error bound tools, the goal is achieving the smallest possible sound error bound, unlike in this paper, where the focus is primarily on computational efficiency.
@@ 9. Conclusion
Evaluating real expressions to high precision is a critical component of algorithms across science, computing, and especially numerics. But state-of-the-art tools use a uniform precision, doubled with every iteration. This wastefully allocates too much precision to non-accuracy-critical operations, and can also overestimate precision or require too many iterations to find the right precision. We instead introduce a fast, sound method for assigning precisions to interval operators so as to correctly-round a real expression. Our method is based on a variation on error Taylor series and condition numbers, called amplification factors, which is sound, computable quickly, and specialized to interval arithmetic. A key optimization is the use of “exponent tricks”, which allow computing amplification factors using only operations on machine integers. As a result, Reval is able to evaluate real expressions 1.47×1.47$\times$1.47 × faster than the state-of-the-art Sollya tool across our suite of 493 benchmarks, with Reval’s advantage rising to 4.92×4.92$\times$4.92 × for the most difficult input points. It achieves this speedup largely by using a sophisticated mixed-precision assignment algorithm that evaluates operations at high precision only when necessary.